OpenStudy (anonymous):

Shaylee has $20,000 property damage insurance, and $5,000 comprehensive insurance with a $500 deductible. One evening, she slides on a a patch of ice, and hits a parked car, causing $4,500 worth of damage to the other car. The accident causes $1,750 of damage to her own car. How much will her insutance company pay to fix both cars? A. 6, 250 B. 5, 750 C. 8, 250 D. 24, 500

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

What is the total cost of repairing all the damage?

OpenStudy (anonymous):

$6,250?

OpenStudy (skullpatrol):

Yes 4500 + 1750 = 6250

OpenStudy (anonymous):

Ah, I see. Thank you. (:

OpenStudy (skullpatrol):

Wait, answer the question...

OpenStudy (skullpatrol):

How much will her insurance company pay to fix both cars?

OpenStudy (anonymous):

I am not sure. It is B? Because 6,250-500= 5, 750?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

Yay. :3

OpenStudy (skullpatrol):

Thanks for trying to understand :)

OpenStudy (anonymous):

I think so. 1750+(4500-500)=5750